An entity has the following cost components for 150,000 units of product for the year:                                                                  

Direct Materials                                                325,000                                                               

Direct Labor                                                       175,000                                                               

Manufacturing Overhead                             225,000                                                              

Selling and Administrative expense          175,000                                                              

                                                                               

All costs are variable except for 75,000 of manufacturing overhead and 75,000 of selling and administrative expenses. The total costs to produce and sell 175,000 units for the year are:                             

Answer:
Answer #1
Variable cost per unit
Direct Materials 2.167 =325000/150000
Direct Labor 1.167
Manufacturing Overhead 1.000 =(225000-75000)/150000
Selling and Administrative expense   0.667 =(175000-75000)/150000
Total 5.000
Total variable costs 875000 =175000*5
Total fixed costs 150000 =75000+75000
Total costs to produce and sell 175,000 units 1025000